OpenStudy (anonymous):

differentiate 5x^3 + 3x^2y + 4y^3

OpenStudy (anonymous):

d/dx(5 x^3+3 x^2 y+4 y^3) Differentiate the sum term by term and factor out constants: = 5 (d/dx(x^3))+3 y (d/dx(x^2))+4 (d/dx(y^3)) The derivative of x^2 is 2 x: = 5 (d/dx(x^3))+4 (d/dx(y^3))+3 (2 x) y The derivative of x^3 is 3 x^2: = 4 (d/dx(y^3))+5 (3 x^2)+6 x y The derivative of y^3 is zero: = 15 x^2+6 x y+4 0

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

np, in return give me a "Good Answer"...;P